# SynHLMA: Synthesizing Hand Language Manipulation for Articulated Object with Discrete Human Object Interaction Representation

### v1

@@ 4.1 Experimental Settings
Implementation Details We use four codebooks with a capacity of 1024 each for the hand and the object. For the large language model, we adopt Vicuna-7B. The model is trained using the LoRA technique, with a rank value set to 16, meaning about 5% of the parameters are involved in the training process. We set the learning rate to 3e-4 and the alpha value to 64, and employ a cosine annealing learning rate scheduler to ensure training stability. The batch size is 128, and the training is conducted for 20 epochs on four RTX A800 GPUs.
Metrics To comprehensively evaluate the quality of the generated grasping instructions from multiple perspectives, we adopt the evaluation methodology proposed by Jiang et al. (2023); Huang et al. (2025). We consider the following evaluation metrics: Fréchet Inception Distance (FID), Diversity, Multi-modal Distance (MMDist), Interaction Volume (IV), Average Displacement Error (ADE), Final Displacement Error (FDE), Codebook Update Coverage (CUC). For detailed definitions, please refer to the supplementary material.
Baselines HAOI generation refers to predicting the subsequent grasping sequence given the object’s point cloud and a corresponding textual description as input. HAOI prediction refers to the task where only the first 20% of a grasping sequence is provided to the model, which must then predict the remaining 80%. HAOI interpolation denotes the scenario in which 40–50% of a sequence is intentionally omitted, requiring the model to complete the missing portions in a coherent manner. We compare SynHLMA with state-of-the-art HAOI generation models, including HOIGPT and Text2HOI, as well as with representative human motion generation baselines such as T2MGPT, MotionGPT, and TM2M.
4.2 Comparison with State-of-the-arts
As shown in Table 1, our method achieves state-of-the-art performance on the HAOI generation task, outperforming all baselines. Notably, it improves the FID score by 4.919% and achieves 13.986% increase in diversity.
Table 2 presents results on the HAOI prediction and interpolation tasks. For HAOI prediction, our method achieves a 14.64% improvement in FID and outperforms baselines by 19.572% in diversity. In the interpolation task, we observe a 9.731% reduction in FID, along with a 19.969% gain in diversity.
Unlike prior models, our method introduces an articulation-guided decoder to explicitly capture articulated object structure, and employs a hierarchical HAOI token representation for fine-grained manipulation modeling. Finally, the proposed articulation-aware loss guides the model to generate grasps that are more consistent with the articulated object’s physical state and kinematic configuration. Qualitative results of the proposed method are illustrated in Figure 5.
@@ 4.3 Ablation Studies
Token Semantics in Discrete Representations We evaluate the impact of token semantics through a series of ablation studies on our stage-wise discrete token design (Table 4). 1) and 2), which remove the object joint token $\langle j\rangle$, result in a significant performance drop and render the refinement token $\langle r\rangle$ ineffective. 3) and 4) introduce additional fine-grained refinement tokens; although they offer slight improvements, they increase training complexity and hinder convergence. In 5), removing all semantic meanings from tokens leads to degraded performance, while 6), which shares a single codebook across all stages, fails to capture the complexity of grasping behaviors.
@@
Table 4: Ablation on VQ-VAE design.
[…]
$\mathcal{B}$ entries $\mathcal{K}$ = 512 1.237 2.944 3.690 0.232
$\mathcal{B}$ entries $\mathcal{K}$ = 2048 0.913 3.063 3.222 0.195
$\mathcal{B}$ dim. $d_{\mathcal{B}}$ = 256 1.113 3.704 3.446 -
$\mathcal{B}$ dim. $d_{\mathcal{B}}$ = 1024 1.055 3.235 3.195 -
EMA + Reset 3.637 3.327 3.704
Multi-Level VQ-VAE Ablations We conduct an ablation study on our multi-level VQ-VAE design (Table 4). Experiments 1 through 4 vary the capacity of the codebook and embedding dimensionality, showing that larger capacities and higher dimensions preserve richer representations and consistently improve all metrics. In 5), we compare training strategies for mitigating codebook gradient collapse. While EMA combined with codebook reset has been reported to help, we find EMA yields limited gains for HAOI generation; thus, our main experiments adopt reset alone. Finally, we track the CUC across evaluations by averaging the three codebooks. Notably, a codebook capacity of only 512 achieves a CUC of 23.2%, indicating that even relatively compact discrete representations suffice for our task.
@@
Methods FID $\downarrow$ MMDist $\downarrow$ IV $\downarrow$
@@
Manipulation Language Model Ablations Our ablation studies on the HAOI manipulation language model are detailed in Table 5. 1) Comparing Llama-7B with instruction-tuned Vicuna-7B reveals the latter’s superior adaptability for text-to-manipulation generation. 2) & 3) Investigating LoRA rank demonstrates that excessively high ranks induce catastrophic forgetting of prior knowledge by increasing trainable parameters; conversely, overly low ranks impede convergence due to underfitting, leading us to select rank 16 as optimal. 4) Compared to two-stage training, the single-stage approach underperforms and exhibits significantly greater instability.
@@ 4.4 Dexterous Manipulation Transfer to Robotics
We demonstrate the application of our SynHLMA in embodied robotics area, and transfer the learned manipulation sequence poses into ShadowHand model Li et al. (2019) within RaiSim simulator in Fig. 6. In detail, we align the finger keypoints from MANO model generated by SynHLMA and ShadowHand model with a simple fitting-based optimization approach. Next, using the fitted ShadowHand poses, we can enable the dexterous manipulation that implements on the same articulated objects.
@@ 5 Conclusion
In this work, we propose SynHLMA, a novel manipulation language model for articulated objects based on discrete representations. To support this framework, we introduce HAOI-Lang, a new dataset that bridges the gap between manipulation actions and their corresponding language instructions. By designing multi-level tokens, our model captures fine-grained grasping processes, achieving strong performance across various downstream tasks. Moreover, a joint-aware loss is incorporated to ensure consistency between the generated hand motions and the articulated object joints. Furthermore, we demonstrate its potential in enhancing dexterous grasping for robotic hands. In future work, we aim to explore more fine-grained and coordinated bimanual manipulation.
Appendix A Appendix
A.1 Overview
In this supplementary material, we first describe in detail the usage of Large Language Models (LLMs) in our paper. We then provide a comprehensive explanation of the composition and generation method of the HAOI-Lan dataset. Next, we present an in-depth description of the experimental setup, including batch loss functions and scoring metrics not covered in the main paper. Finally, we showcase the visualization results of our Discrete Articulated Manipulation Representation across multiple downstream HAOI tasks, demonstrating its generalization ability across different objects within the same category.
A.2 Use of Large Language Models
@@
A.3 HAOI-Lang Dataset Construction
A.3.1 Dataset Composition and Object Coverage
Our HAOI-Lan dataset consists of seven commonly encountered articulated objects: stapler, laptop, scissors, cabinet, dishwasher, eyeglasses, and box. These objects collectively cover nearly all major types of articulated structures found in real-world environments, such as rotational joints, sliding mechanisms, and compound linkages. Each object category includes up to 50 diverse instances, ensuring significant intra-class variation in appearance, articulation, and geometry.
@@ A.3.1 Dataset Composition and Object Coverage
Category Instances Angle Manipulations & Captions Static Grasp
Stapler 8 10 1600 16000
Box 8 10 1600 16000
Laptop 55 10 11000 110000
Scissors 45 10 9000 90000
Cabinet 38 10 7600 76000
Dishwasher 37 10 7400 74000
Eyeglasses 65 10 13000 130000
Total 256 - 51200 512000
A.3.2 Semantic Captioning and Instructional Dialogue Generation
@@
A.4 Experiments Details
A.4.1 Implementation Details
@@
  • $\lambda_{P}$ : 1.0
  • $\lambda_{C}$ : 1.0
  • $\lambda_{J}$ : 1.0
@@
A.4.2 Evaluation Metrics
To complement the evaluation methodology discussed in the main paper, we provide a detailed account of the quantitative metrics used to assess the quality, diversity, and accuracy of the generated hand-object interaction sequences. These metrics are inspired by previous work in instruction-conditioned motion generation Jiang et al. (2023); Huang et al. (2025) and are tailored to capture both spatial precision and behavioral richness. In addition, to verify that the capacity of the VQ-VAE can adequately discretize continuous tasks, we designed a metric to measure its utilization rate.
@@ A.4.2 Evaluation Metrics
A.5 Generalization Across Object Scales in HAOI Tasks
@@
The visualizations (see Figure 8 and Figure 9,10) show that:
@@
Figure 8: HAOI generation, prediction, and completion on small-to-medium objects: laptop and glasses.

### v2

@@ 4.1 Experimental Settings
Implementation and metrics We employ four codebooks to discretize hand and object representations. The manipulation language model is built upon Vicuna-7B-v1.5 and fine-tuned using LoRA. To comprehensively evaluate the quality of the generated grasping instructions from multiple perspectives, we adopt the evaluation methodology proposed by Jiang et al. (2023); Huang et al. (2025). We consider the following evaluation metrics: Fréchet Inception Distance (FID), Diversity, Multi-modal Distance (MMDist), Interaction Volume (IV), Average Displacement Error (ADE), Final Displacement Error (FDE), Codebook Update Coverage (CUC). For detailed configurations and definitions, please refer to the supplementary material.
@@
Baselines HAOI generation refers to predicting the subsequent grasping sequence given the object’s point cloud and a corresponding textual description as input. HAOI prediction refers to the task where only the first 20% of a grasping sequence is provided to the model, which must then predict the remaining 80%. HAOI interpolation denotes the scenario in which 40–50% of a sequence is intentionally omitted, requiring the model to complete the missing portions in a coherent manner. We compare SynHLMA with state-of-the-art HAOI generation models, including HOIGPT, Text2HOI, SemGrasp and NL2Contact, as well as with representative human motion generation baselines such as T2MGPT, MotionGPT, and TM2M.
4.2 Comparison with State-of-the-arts
As shown in Table 1, our method achieves state-of-the-art performance on the HAOI generation task, outperforming all baselines. Notably, it improves the FID score by 4.919% and achieves 12.530% increase in diversity. Table 2 presents results on the HAOI prediction and interpolation tasks. For HAOI prediction, our method achieves a 14.64% improvement in FID and outperforms baselines by 19.572% in diversity. In the interpolation task, we observe a 9.731% reduction in FID, along with a 19.969% gain in diversity. Unlike prior models, our method introduces an articulation-guided decoder to explicitly capture articulated object structure, and employs a hierarchical HAOI token representation for fine-grained manipulation modeling. Finally, the proposed articulation-aware objective guides the model to generate grasps that are more consistent with the articulated object’s physical state and kinematic configuration. Qualitative results of the proposed method are illustrated in Figure 5.
@@ 4.3 Ablation Studies
Articulation-Aware Objective Ablations Table 3 reports the quantitative results of the articulation-aware objective. Removing $\mathcal{L}_{geom}$ increases FID to 15.872 and IV to 6.452, while removing $\mathcal{L}_{temp}$ raises ADE to 1.112; ablating both leads to the worst performance with FID 16.840, IV 6.987, and ADE 1.233. These results demonstrate that our articulation-aware objective effectively enforces geometric validity and temporal consistency, enabling joint-state-aligned motion generation that faithfully follows articulated object dynamics.
Multi-Level VQ-VAE Ablations Table 4 presents the ablation results on the multi-level VQ-VAE design. Increasing the codebook size to $\mathcal{K}=2048$ achieves the best FID of 0.913 and Diversity of 3.063, while setting $d_{\mathcal{B}}=1024$ yields the lowest IV of 3.195; even a compact codebook with $\mathcal{K}=512$ attains a CUC of 0.232. This indicates that our hierarchical discrete representation preserves rich motion semantics while maintaining compactness, validating the effectiveness of structured codebook design for articulated grasp generation.
Token Semantics in Discrete Representations Ablations Table 5 summarizes the ablation results on token semantics. Our full design $\langle g,l,r,j\rangle$ achieves the best FID of 0.699 and ADE of 0.815, outperforming variants such as removing all tokens or discarding semantic meanings. These findings confirm that explicitly encoding stage-wise semantic roles enables structured motion decomposition and enhances the model’s capacity to capture fine-grained grasping behaviors.
Manipulation Language Model Ablations Table 6 reports the ablation results on the manipulation language model. Among different backbones, Gemma achieves FID 22.576 and IV 2.317, while Llama yields FID 51.911; inappropriate LoRA ranks severely degrade performance to FID 126.954, and removing the two-stage training results in FID 39.849. These results indicate that a properly configured backbone together with staged optimization stabilizes language-to-motion alignment, facilitating accurate mapping from textual manipulation intent to physically grounded action sequences.
Table 3: Ablation on articulation-aware objective.
FID $\downarrow$ IV $\downarrow$ ADE $\downarrow$
w/o $\mathcal{L}_{geom}$ 15.872 6.452 1.041
w/o $\mathcal{L}_{temp}$ 14.990 6.104 1.112
w/o $\mathcal{L}_{geom}\land\mathcal{L}_{temp}$ 16.840 6.987 1.233
Table 4: Ablation on VQ-VAE design.
FID $\downarrow$ Diversity $\uparrow$ IV $\downarrow$ CUC
$\mathcal{B}$ entries $\mathcal{K}$ = 512 1.237 2.944 3.690 0.232
$\mathcal{B}$ entries $\mathcal{K}$ = 2048 0.913 3.063 3.222 0.195
$\mathcal{B}$ dim. $d_{\mathcal{B}}$ = 256 1.113 3.704 3.446 -
$\mathcal{B}$ dim. $d_{\mathcal{B}}$ = 1024 1.055 3.235 3.195 -
EMA + Reset 3.637 3.327 3.704 -
@@
FID $\downarrow$ MMDist $\downarrow$ IV $\downarrow$
@@ 4.4 Dexterous Manipulation Transfer to Robotics
We demonstrate the application of our SynHLMA in embodied robotics area, and transfer the learned manipulation sequence poses into ShadowHand model (Li et al., 2019) within RaiSim simulator in Fig. 6. In detail, we align the finger keypoints from MANO model generated by SynHLMA and ShadowHand model with a simple fitting-based optimization approach. Next, using the fitted ShadowHand poses, we can enable the dexterous manipulation that implements on the same articulated objects.
@@ 5 Conclusion
In this work, we propose SynHLMA, a novel manipulation language model for articulated objects based on discrete representations. Moreover, an articulation-aware objective is incorporated to enforce geometric and physical validity, achieve representation-level joint-state alignment, and enhance the model’s responsiveness to articulated object variations. To support this framework, we introduce HAOI-Lang, a new dataset that bridges the gap between manipulation actions and their corresponding language instructions. By designing multi-level tokens, our model captures fine-grained grasping processes, achieving strong performance across various downstream tasks. Furthermore, we demonstrate its potential in enhancing dexterous grasping for robotic hands. In future work, we aim to explore more fine-grained and coordinated bimanual manipulation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Overview
In this supplementary material, we first describe in detail the usage of Large Language Models (LLMs) in our paper. We then provide a comprehensive explanation of the composition and generation method of the HAOI-Lang dataset. Next, we present an in-depth description of the experimental setup, including batch loss functions and scoring metrics not covered in the main paper. Finally, we showcase the visualization results of our Discrete Articulated Manipulation Representation across multiple downstream HAOI tasks, demonstrating its generalization ability across different objects within the same category.
Appendix B Use of Large Language Models
@@
Appendix C HAOI-Lang Dataset Construction
C.1 Dataset Composition and Object Coverage
Our HAOI-Lan dataset consists of seven commonly encountered articulated objects: stapler, laptop, scissors, cabinet, dishwasher, eyeglasses, and box. These objects collectively cover nearly all major types of articulated structures found in real-world environments, such as rotational joints, sliding mechanisms, and compound linkages. Each object category includes around 30–70 object instances, ensuring significant intra-class variation in appearance, articulation, and geometry.
@@ C.1 Dataset Composition and Object Coverage
Category Instances Angle Manipulations & Captions
Stapler 8 10 1600
Box 8 10 1600
Laptop 55 10 11000
Scissors 45 10 9000
Cabinet 38 10 7600
Dishwasher 37 10 7400
Eyeglasses 65 10 13000
Total 256 - 51200
C.2 Semantic Captioning and Instructional Dialogue Generation
@@
Figure 7: Instructional dialogue template created by GPT-4 for HAOI-Lang.
Appendix D Experiments Details
D.1 Implementation Details
@@
  • $\lambda_{\text{pen}}$ : 1.0
  • $\lambda_{\text{joint}}$ : 1.0
  • $\lambda_{\text{temp}}$ : 1.0
@@
D.2 Evaluation Metrics
To complement the evaluation methodology discussed in the main paper, we provide a detailed account of the quantitative metrics used to assess the quality, diversity, and accuracy of the generated hand-object interaction sequences. These metrics are inspired by previous work in instruction-conditioned motion generation (Jiang et al., 2023; Huang et al., 2025) and are tailored to capture both spatial precision and behavioral richness. In addition, to verify that the capacity of the VQ-VAE can adequately discretize continuous tasks, we designed a metric to measure its utilization rate.
@@ D.2 Evaluation Metrics
Appendix E Generalization Across Object Scales in HAOI Tasks
@@
The visualizations (see Figure 8, Figure 9, Figure 10 and 11) show that:
@@
Appendix F Failure Cases in HAOI Manipulation
We select the most representative failure cases from the test set, as illustrated in Fig. 12. In HAOI generation, when the manipulation description is overly strict or physically implausible, the resulting grasp may deviate from realistic or physically feasible configurations. In HAOI prediction, the predicted interaction sequence may exhibit temporal inconsistencies, such as hand drift from the intended grasping position or deformation of the grasp over time. In HAOI interpolation, the primary issue is abrupt changes in hand pose, which are often caused by discontinuities or invalid predictions in the interpolated pose sequence.
Figure 8: HAOI generation, prediction, and completion on small-to-medium objects: laptop.
Figure 9: HAOI generation, prediction, and completion on small-to-medium objects: glasses.